Required information

[The following information applies to the questions displayed below.]

Pacific Ink had beginning work-in-process inventory of $1,005,960 on October 1. Of this amount, $423,500 was the cost of direct materials and $582,460 was the cost of conversion. The 61,000 units in the beginning inventory were 30 percent complete with respect to both direct materials and conversion costs.

During October, 128,000 units were transferred out and 43,000 remained in ending inventory. The units in ending inventory were 80 percent complete with respect to direct materials and 40 percent complete with respect to conversion costs. Costs incurred during the period amounted to $3,314,300 for direct materials and $4,022,730 for conversion.

Required:

a. Compute the equivalent units for the materials and conversion cost calculations.

b. Compute the cost per equivalent unit for direct materials and for conversion costs using the FIFO method.

c. Compute the cost of goods transferred out and the cost of ending inventory using the FIFO method.

d. Is the ending inventory higher or lower under the weighted-average method compared to FIFO?
Answer #1

a) Equivalent unit

Whole Unit % material EUP-material % conversion EUP-conversion
Beginning WIP 61000 70% 42700 70% 42700
Unit started and completed 67000 100% 67000 100% 67000
Ending WIP 43000 80% 34400 40% 17200
Equivalent unit 171000 144100 126900

b) Cost per equivalent unit

Material Conversion
Cost added 3314300 4022730
Equivalent unit 144100 126900
Cost per equivalent unit 23 31.70

c) Cost of goods transferred out and ending inventory

Direct material Conversion Total
Cost of unit transferred out (109700*23+423500) = 2946600 (109700*31.70+582460) = 4059950 7006550
Ending work in process 34400*23 = 791200 17200*31.70 = 545240 1336440

d) Ending inventory would be higher weighted average method